# Deep Reinforcement Learning for Infinite Horizon Mean Field Problems in Continuous Spaces

### v1

@@ 5.4 Hyperparameters and Numerical Specifics
Figure 4: The blue curve is a rolling average of the absolute error between the mean of samples produced from the parameterized score function $\Sigma_{\varphi_{n}}$ and the optimal mean $\hat{m}$ from eq. 5.5 in the case of MFG (left) and $m^{*}$ from eq. 5.8 in the case of MFC (right). Large jumps are due to random outliers which result from the stochasticity of our algorithm.
@@ 6 Actor-Critic Algorithm for Mean Field Control Games (MFCG)
As observed in [Angiuli et al., 2023a] in the case of tabular Q-learning, our IH-MF-AC algorithm (Algorithm 1) can easily be extended to the case of mixed mean field control game problems that involve two population distributions, a local one and a global one. This type of game corresponds to competitive games between a large number of large collaborative groups of agents. The local distribution is the “representative" agent’s group distribution, while the global distribution is the distribution of the entire population. We refer to [Angiuli et al., 2023a, Angiuli et al., 2022a] for further details on MFCG, including the limit from finite player games to infinite player games. Note that the solution gives an approximation of the Nash equilibrium between the competitive groups.
@@
0:  Initial distribution $\xi$; number of time steps $N\gg 0$; discrete time step size $\Delta t$; neural network learning rates for actor $\rho_{\Pi}$, critic $\rho_{V}$, global score $\rho_{\Sigma}$, and local score $\rho_{\widetilde{\Sigma}}$; Langevin dynamics step size $\epsilon$.
@@
2:  Sample $X_{t_{0}}\sim\xi$
@@
4:     Compute score loss for $\Sigma$: $\quad L_{\Sigma}(\varphi_{n})=\operatorname{tr}\left(\nabla_{x}\Sigma_{\varphi_{n}}(X_{t_{n}})\right)+\frac{1}{2}\left\lVert\Sigma_{\varphi_{n}}(X_{t_{n}})\right\rVert_{2}^{2}$
5:     Update $\Sigma$ with SGD: $\quad\varphi_{n+1}=\varphi_{n}-\rho_{\Sigma}\nabla_{\varphi}L_{\Sigma}(\varphi_{n})$
6:     Compute score loss for $\widetilde{\Sigma}$: $\quad L_{\widetilde{\Sigma}}(\xi_{n})=\operatorname{tr}\left(\nabla_{x}\widetilde{\Sigma}_{\xi_{n}}(X_{t_{n}})\right)+\frac{1}{2}\left\lVert\widetilde{\Sigma}_{\xi_{n}}(X_{t_{n}})\right\rVert_{2}^{2}$
7:     Update $\widetilde{\Sigma}$ with SGD: $\quad\xi_{n+1}=\xi_{n}-\rho_{\widetilde{\Sigma}}\nabla_{\xi}L_{\widetilde{\Sigma}}(\xi_{n})$
8:     Generate mean field samples $S_{t_{n}}=\left(S_{t_{n}}^{(1)},S_{t_{n}}^{(2)},\dots,S_{t_{n}}^{(k)}\right)$ from $\Sigma_{\varphi_{n+1}}$ and $\widetilde{S}_{t_{n}}=\left(\widetilde{S}_{t_{n}}^{(1)},\widetilde{S}_{t_{n}}^{(2)},\dots,\widetilde{S}_{t_{n}}^{(k)}\right)$ from $\widetilde{\Sigma}_{\xi_{n+1}}$ using Langevin dynamics (eq. 4.3) with step size $\epsilon$ and compute $\overline{\mu}_{S_{t_{n}}}\coloneqq\frac{1}{k}\sum_{i=1}^{k}\delta_{S_{t_{n}}^{(i)}}$ and $\overline{\mu}_{\widetilde{S}_{t_{n}}}\coloneqq\frac{1}{k}\sum_{i=1}^{k}\delta_{\widetilde{S}_{t_{n}}^{(i)}}$.
9:     Sample action: $\quad A_{t_{n}}\sim\Pi_{\psi_{n}}(\cdot\mid X_{t_{n}})$
10:     Observe reward from the environment: $\quad r_{n+1}=-f(X_{t_{n}},\overline{\mu}_{S_{t_{n}}},\overline{\mu}_{\widetilde{S}_{t_{n}}}A_{t_{n}})\Delta t$
11:     Observe next state from environment:$\quad X_{t_{n+1}}=b(X_{t_{n}},\overline{\mu}_{S_{t_{n}}},A_{t_{n}})\Delta t+\sigma(X_{t_{n}},\overline{\mu}_{S_{t_{n}}},A_{t_{n}})\sqrt{\Delta t}\,z_{n},\qquad z_{n}\sim\mathcal{N}(0,1)$
12:     Compute TD target: $\quad y_{n+1}=r_{n+1}+e^{-\beta\Delta t}V_{\theta_{n}}(X_{t_{n+1}})$
13:     Compute TD error: $\quad\delta_{\theta_{n}}=y_{n+1}-V_{\theta_{n}}(X_{t_{n}})$
14:     Compute critic loss: $\quad L_{V}(\theta_{n})=\delta_{\theta_{n}}^{2}$
15:     Update critic with SGD: $\quad\theta_{n+1}=\theta_{n}-\rho_{V}\nabla_{\theta}L_{V}(\theta_{n})$
16:     Compute actor loss: $\quad L_{\Pi}(\psi_{n})=-\delta_{\theta_{n}}\log\Pi_{\psi_{n}}(A_{t_{n}}\mid X_{t_{n}})$
17:     Update actor with SGD: $\quad\psi_{n+1}=\psi_{n}-\rho_{\Pi}\nabla_{\psi}L_{\Pi}(\psi_{n})$
18:  end for
19:  return  $(\Pi_{\psi_{N}},\Sigma_{\varphi_{N}},\widetilde{\Sigma}_{\xi_{N}})$
@@ 6.1 A Linear-Quadratic Benchmark
We present the analytic solution to the MFCG problem using notation consistent with the derivation in [Angiuli et al., 2023a]. Define
$\Gamma_{2}=\frac{-\beta+\sqrt{\beta^{2}+8\left(c_{1}+c_{3}+\tilde{c}_{1}\right)}}{4}\qquad\text{and}\qquad\Gamma_{1}=-\frac{2\Gamma_{2}c_{3}c_{4}}{c_{1}\left(1-c_{2}\right)+\tilde{c}_{1}\left(1-\tilde{c}_{2}\right)^{2}+c_{3}+\tilde{c}_{5}}.$
@@ 6.2 Hyperparameters and Numerical Specifics
The results of the IH-MFCG-AC algorithm (Algorithm 2) are presented in figs. 6 and 6. As expected, the learning of the global and local distributions reflects that of the optimal MFG distribution and the optimal MFC distribution, respectively. We observe that the global score is learned faster and with more accuracy than the local score, which is prone to outliers and instability. The optimal control is learned well within the support of the optimal distribution, but could possibly be expanded with a more advanced exploration strategy.
Figure 5: The histograms are the learned distributions generated using samples from the global score $\Sigma_{\varphi_{n}}$ (green) representing the global distribution $\hat{\mu}$ and the local score $\widetilde{\Sigma}_{\xi_{n}}$ (purple) representing the local score $\mu^{\alpha,\hat{\mu}}$ after $N=2\times 10^{6}$ iterations. The dashed line (blue) is the learned feedback control. The benchmark solution to the MFCG is provided in orange. The $x$-axis shows the state variable $x$, the left $y$-axis refers to the value of the control $\alpha(x)$, and the right axis represents the probability density of $\mu(x)$.
Figure 5: The histograms are the learned distributions generated using samples from the global score $\Sigma_{\varphi_{n}}$ (green) representing the global distribution $\hat{\mu}$ and the local score $\widetilde{\Sigma}_{\xi_{n}}$ (purple) representing the local score $\mu^{\alpha,\hat{\mu}}$ after $N=2\times 10^{6}$ iterations. The dashed line (blue) is the learned feedback control. The benchmark solution to the MFCG is provided in orange. The $x$-axis shows the state variable $x$, the left $y$-axis refers to the value of the control $\alpha(x)$, and the right axis represents the probability density of $\mu(x)$.
Figure 6: The blue curve is a rolling average of the absolute error of the mean of samples produced from the global score function $\Sigma_{\varphi_{n}}$ (left)—denoted $\hat{m}_{t_{n}}$—and the local score function $\widetilde{\Sigma}_{\xi_{n}}$—denoted $m^{\alpha,\hat{\mu}}_{t_{n}}$—compared to the optimal mean $m$ from eq. 6.8. Large jumps are due to random outliers which result from the stochasticity of our algorithm.

### v3

@@ 5.5 Hyperparameters and Numerical Specifics
Figure 2: We plot the absolute error between the mean of samples produced from the parameterized score function $\Sigma_{\varphi_{n}}$ and the optimal mean $\hat{m}$ in the case of MFG (left) and $m^{*}$ in the case of MFC (right). These values were averaged over five runs each with different random initial samples with the standard deviation given by the light blue shaded region. Large jumps are due to random outliers which result from the stochasticity of our algorithm.
(a) Results for $\rho_{\Sigma}=10^{-6}$
(b) Results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 3: The orange and green curves are the optimal value functions for the MFG and MFC problem, respectively. The blue dashed line is the learned value function given by the negative of critic $V_{\theta_{N}}$ averaged over five runs with different initial samples after $N=10^{6}$ iterations. Since the original optimization problem aims to minimize cost while our algorithm seeks to maximize reward, we take the negative of the critic function to make the problems equivalent. The light blue shaded region depicts one standard deviation from the learned value.
(c) Results for $\rho_{\Sigma}=10^{-6}$
(d) Results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 4: We plot the absolute error between the expected error between the learned value function $V_{\theta_{n}}$ and the optimal value function $\hat{v}$ in the case of MFG (left) and $v^{*}$ in the case of MFC (right). These plots were averaged over five runs each with different random initial samples and with the standard deviation given by the light blue shaded region. Large jumps are due to random outliers which result from the stochasticity of our algorithm.
(a) Results for $\rho_{\Sigma}=10^{-6}$
(b) Results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 5: We plot the expected error between the learned control function $\alpha_{n}=\mathbb{E}[\Pi_{\psi_{n}}]$ and the optimal control $\hat{\alpha}$ in the case of MFG (left) and $\alpha^{*}$ in the case of MFC (right). These plots were averaged over five runs each with different random initial samples and with the standard deviation given by the light blue shaded region. Large jumps are due to random outliers which result from the stochasticity of our algorithm.
Table 4: Running cost coefficients and volatility for eqs. 5.1 and 5.2. The results for this parameter set are displayed in figs. 7, 7, 9, 9 and 10.
$c_{1}$ $c_{2}$ $c_{3}$ $c_{4}$ $c_{5}$ $\sigma$
0.15 1.0 0.25 1.0 2.0 0.5
(a) MFG results for $\rho_{\Sigma}=10^{-6}$
(b) MFC results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 6: The histogram (blue) is the learned asymptotic distribution using samples generated from $\Sigma_{\varphi_{n}}$ and the dashed line (blue) is the learned feedback control after $N=10^{6}$ iterations averaged over five runs with different initial samples. The green curves correspond to the optimal control and mean field distribution for MFC, while the orange curves are the equivalent for MFG. The bottom axis shows the state variable $x$, the left axis refers to the value of the control $\alpha(x)$, and the right axis represents the probability density of $\mu$.
(c) Results for $\rho_{\Sigma}=10^{-6}$
(d) Results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 7: We plot the absolute error between the mean of samples produced from the parameterized score function $\Sigma_{\varphi_{n}}$ and the optimal mean $\hat{m}$ in the case of MFG (left) and $m^{*}$ in the case of MFC (right). These values were averaged over five runs each with different random initial samples with the standard deviation given by the light blue shaded region. Large jumps are due to random outliers which result from the stochasticity of our algorithm.
(a) Results for $\rho_{\Sigma}=10^{-6}$
(b) Results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 8: The orange and green curves are the optimal value functions for the MFG and MFC problem, respectively. The blue dashed line is the learned value function given by the negative of critic $V_{\theta_{N}}$ averaged over five runs with different initial samples after $N=10^{6}$ iterations. Since the original optimization problem aims to minimize cost while our algorithm seeks to maximize reward, we take the negative of the critic function to make the problems equivalent. The light blue shaded region depicts one standard deviation from the learned value.
(c) Results for $\rho_{\Sigma}=10^{-6}$
(d) Results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 9: We plot the absolute error between the expected error between the learned value function $V_{\theta_{n}}$ and the optimal value function $\hat{v}$ in the case of MFG (left) and $v^{*}$ in the case of MFC (right). These plots were averaged over five runs each with different random initial samples and with the standard deviation given by the light blue shaded region. Large jumps are due to random outliers which result from the stochasticity of our algorithm.
(a) Results for $\rho_{\Sigma}=10^{-6}$
(b) Results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 10: We plot the expected error between the learned control function $\alpha_{n}=\mathbb{E}[\Pi_{\psi_{n}}]$ and the optimal control $\hat{\alpha}$ in the case of MFG (left) and $\alpha^{*}$ in the case of MFC (right). These plots were averaged over five runs each with different random initial samples and with the standard deviation given by the light blue shaded region. Large jumps are due to random outliers which result from the stochasticity of our algorithm.
$\sigma$ 0.3 0.5 0.7 0.9 1.1
$e_{\hat{m}}$ 0.087 0.073 0.284 0.591 0.981
$e_{\hat{\alpha}}$ 0.267 0.242 0.693 0.802 1.819
$e_{\hat{v}}$ 0.178 0.394 0.522 0.673 2.005
Table 5: Error metrics for $\rho_{\Sigma}=10^{-6}$. For increasing values of $\sigma$, we list the $L^{2}$ error in the learned mean $e_{\hat{m}}$, the expected $L^{2}$ error in the learned control $e_{\hat{\alpha}}$, and the expected $L^{2}$ error in the learned value function $e_{\hat{v}}$ after $N=10^{6}$ times steps.
$\sigma$ 0.3 0.5 0.7 0.9 1.1
$e_{m^{*}}$ 0.115 0.299 1.876 2.701 3.415
$e_{\alpha^{*}}$ 0.583 0.441 1.106 4.556 6.946
$e_{v^{*}}$ 0.455 0.742 2.342 6.035 5.569
Table 6: Error metrics for $\rho_{\Sigma}=5\times 10^{-4}$. For increasing values of $\sigma$, we list the $L^{2}$ error in the learned mean $e_{m^{*}}$, the expected $L^{2}$ error in the learned control $e_{\alpha^{*}}$, and the expected $L^{2}$ error in the learned value function $e_{v^{*}}$ after $N=10^{6}$ time steps.
We have shown that the same algorithm tested on a continuous-state space infinite horizon LQ problem, with hyperparameters shown in Table 1, recovers the MFG or MFC solution. However, our numerical experiments show that there is more stability for the MFG problem, as also observed in the case of the Q-learning algorithm in the context of infinite horizon and finite-space [Angiuli et al., 2022b], and in [Angiuli et al., 2023b] in the context of discrete-space finite horizon problems. In the following section, we see that we observe a similar phenomenon in the case of the mixed mean field control game. Note, however, that in our general algorithm we are not taking advantage of the fact that the numerical example is linear-quadratic. Doing so, we would know a priori that the value function is quadratic and that the control is linear, as is done, for instance, in [Chen et al., 2023].
Table 7: Running cost coefficients and volatility for eqs. 5.9 and 5.10. The results for this parameter set are displayed in figs. 13, 12, 12, 15, 15 and 16. The coefficient values below were randomly generated with the exception of $\sigma$.
$C_{1}$ $C_{2}$ $C_{3}$ $c_{4}$ $C_{5}$ $\sigma$
$\begin{bmatrix}0.964&0.236\\ 0.236&0.076\end{bmatrix}$ $\begin{bmatrix}0.677&0.937\\ 0.937&1.357\end{bmatrix}$ $\begin{bmatrix}0.988&1.188\\ 1.188&1.483\end{bmatrix}$ $\begin{bmatrix}0.810\\ 0.872\end{bmatrix}$ $\begin{bmatrix}0.011&0.072\\ 0.072&0.520\end{bmatrix}$ $\begin{bmatrix}0.3&0\\ 0&0.3\end{bmatrix}$
(a) Results for $\rho_{\Sigma}=10^{-6}$
(b) Results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 11: The scatter plot of points (blue) are the samples of the learned asymptotic distribution generated from the parameterized score function $\Sigma_{\varphi_{n}}$ after $N=10^{6}$ iterations. The solid ellipses are the set of points with Mahalanobis distance three from the optimal mean field distributions in the case of MFG (orange) and MFC (green).
(c) Results for $\rho_{\Sigma}=10^{-6}$
(d) Results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 12: We plot the absolute error between the mean of samples produced from the parameterized score function $\Sigma_{\varphi_{n}}$ and the optimal mean $\hat{m}$ in the case of MFG (left) and $m^{*}$ in the case of MFC (right). These values were averaged over five runs each with different random initial samples with the standard deviation given by the light blue shaded region. Large jumps are due to random outliers which result from the stochasticity of our algorithm.
(a) Results for $\rho_{\Sigma}=10^{-6}$
(b) Results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 13: The orange and green vector fields (right) are the optimal controls for the MFG (top) and MFC (bottom) problems, respectively. The blue vector fields (left) show the learned feedback controls after $N=10^{6}$ iterations averaged over five runs with different initial samples.
(a) Results for $\rho_{\Sigma}=10^{-6}$
(b) Results for $\rho_{\Sigma}=10^{-5}$
Figure 14: We plot the expected error between the learned control function $\alpha_{n}=\mathbb{E}[\Pi_{\psi_{n}}]$ and the optimal control $\hat{\alpha}$ in the case of MFG (left) and $\alpha^{*}$ in the case of MFC (right). These plots were averaged over five runs each with different random initial samples and with the standard deviation given by the light blue shaded region. Large jumps are due to random outliers which result from the stochasticity of our algorithm.
(c) Results for $\rho_{\Sigma}=10^{-6}$
(d) Results for $\rho_{\Sigma}=10^{-5}$
Figure 15: We plot the absolute error between the expected error between the learned value function $V_{\theta_{n}}$ and the optimal value function $\hat{v}$ in the case of MFG (left) and $v^{*}$ in the case of MFC (right). These plots were averaged over five runs each with different random initial samples and with the standard deviation given by the light blue shaded region. Large jumps are due to random outliers which result from the stochasticity of our algorithm.
(a) Results for $\rho_{\Sigma}=10^{-6}$
(b) Results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 16: The right-hand side plots are the optimal value functions for the MFG (top) and MFC (bottom) problems, respectively. The left-hand side plots show the learned functions after $N=10^{6}$ iterations averaged over five runs with different initial samples.
@@ 6 Actor-Critic Algorithm for Mean Field Control Games (MFCG)
As observed in [Angiuli et al., 2023a] in the case of tabular Q-learning, our IH-MF-AC algorithm (Algorithm 1) can easily be extended to the case of mixed mean field control game problems that involve two population distributions, a local one and a global one. This type of game corresponds to competitive games among a large number of large groups, where agents within each group collaborate. The local distribution represents the distribution within each "representative" agent’s group, while the global distribution represents the distribution of the entire population.
Such games are motivated as follows: Consider a finite population of agents consisting of $M$ groups, each containing $N$ agents. An agent indexed by $(m,n)$ indicates that she is the $n^{th}$ member of the $m^{th}$ group. Agents collaborate within their respective groups (sharing the same first index) and compete with all agents from other groups. In other words, all $N$ agents in group $m$ collectively work to minimize the total cost of group $m$. The MFCG corresponds to the mean-field limit as $N$ and $M$ tend to infinity, and serves as a proxy for the solution in the finite player case. We refer to [Angiuli et al., 2023a, Angiuli et al., 2022a] for further details on MFCG, including the limit from finite player games to infinite player games. Note that the solution gives an approximation of the Nash equilibrium between the competitive groups.
@@
0:  Number of time steps $N\gg 0$; discrete time step size $\Delta t$; neural network learning rates for actor $\rho_{\Pi}$, critic $\rho_{V}$, global score $\rho_{\Sigma}$, and local score $\rho_{\widetilde{\Sigma}}$; Langevin dynamics step size $\epsilon$.
@@
2:  Agent receives initial state $X_{t_{0}}$ from the Environment.
@@
4:     Environment computes score loss for $\Sigma$: $\quad L_{\Sigma}(\varphi_{n})=\operatorname{tr}\left(\nabla_{x}\Sigma_{\varphi_{n}}(X_{t_{n}})\right)+\frac{1}{2}\left\lVert\Sigma_{\varphi_{n}}(X_{t_{n}})\right\rVert_{2}^{2}$
5:     Environment updates $\Sigma$ with SGD: $\quad\varphi_{n+1}=\varphi_{n}-\rho_{\Sigma}\nabla_{\varphi}L_{\Sigma}(\varphi_{n})$
6:     Environment computes score loss for $\widetilde{\Sigma}$: $\quad L_{\widetilde{\Sigma}}(\xi_{n})=\operatorname{tr}\left(\nabla_{x}\widetilde{\Sigma}_{\xi_{n}}(X_{t_{n}})\right)+\frac{1}{2}\left\lVert\widetilde{\Sigma}_{\xi_{n}}(X_{t_{n}})\right\rVert_{2}^{2}$
7:     Environment updates $\widetilde{\Sigma}$ with SGD: $\quad\xi_{n+1}=\xi_{n}-\rho_{\widetilde{\Sigma}}\nabla_{\xi}L_{\widetilde{\Sigma}}(\xi_{n})$
8:     Environment generates mean field samples $S_{t_{n}}=\left(S_{t_{n}}^{(1)},S_{t_{n}}^{(2)},\dots,S_{t_{n}}^{(k)}\right)$ from $\Sigma_{\varphi_{n+1}}$ and $\widetilde{S}_{t_{n}}=\left(\widetilde{S}_{t_{n}}^{(1)},\widetilde{S}_{t_{n}}^{(2)},\dots,\widetilde{S}_{t_{n}}^{(k)}\right)$ from $\widetilde{\Sigma}_{\xi_{n+1}}$ using Langevin dynamics (eq. 4.3) with step size $\epsilon$ and compute $\overline{\mu}_{S_{t_{n}}}\coloneqq\frac{1}{k}\sum_{i=1}^{k}\delta_{S_{t_{n}}^{(i)}}$ and $\overline{\mu}_{\widetilde{S}_{t_{n}}}\coloneqq\frac{1}{k}\sum_{i=1}^{k}\delta_{\widetilde{S}_{t_{n}}^{(i)}}$.
9:     Agent samples action: $\quad A_{t_{n}}\sim\Pi_{\psi_{n}}(\cdot\mid X_{t_{n}})$
10:     Agent observes reward $r_{n+1}$ and next state $X_{t_{n+1}}$from the environment.
11:     Agent computes TD target: $\quad y_{n+1}=r_{n+1}+e^{-\beta\Delta t}V_{\theta_{n}}(X_{t_{n+1}})$
12:     Agent computes TD error: $\quad\delta_{\theta_{n}}=y_{n+1}-V_{\theta_{n}}(X_{t_{n}})$
13:     Agent computes critic loss: $\quad L_{V}(\theta_{n})=\delta_{\theta_{n}}^{2}$
14:     Agent updates critic with SGD: $\quad\theta_{n+1}=\theta_{n}-\rho_{V}\nabla_{\theta}L_{V}(\theta_{n})$
15:     Agent computes actor loss: $\quad L_{\Pi}(\psi_{n})=-\delta_{\theta_{n}}\log\Pi_{\psi_{n}}(A_{t_{n}}\mid X_{t_{n}})$
16:     Agent updates actor with SGD: $\quad\psi_{n+1}=\psi_{n}-\rho_{\Pi}\nabla_{\psi}L_{\Pi}(\psi_{n})$
17:  end for
18:  return  $(\Pi_{\psi_{N}},\Sigma_{\varphi_{N}},\widetilde{\Sigma}_{\xi_{N}})$
@@ 6.1 A Linear-Quadratic Benchmark
We present the analytic solution to the MFCG problem using notation consistent with the derivation in [Angiuli et al., 2023a]. The value function is defined as
$\begin{split}v(x)\coloneqq\inf_{\alpha\in\mathbb{A}}\mathbb{E}\Biggl{[}\int_{0}^{\infty}e^{-\beta t}\biggl{(}\frac{1}{2}\alpha_{t}^{2}+c_{1}\left(\mathrm{X}_{t}^{\alpha,\mu}-c_{2}m\right)^{2}+c_{3}\left(\mathrm{X}_{t}^{\alpha,\mu}-c_{4}\right)^{2}\\ \quad{}+\tilde{c}_{1}\left(\mathrm{X}_{t}^{\alpha,\mu}-\tilde{c}_{2}m^{\alpha,\mu}\right)^{2}+\tilde{c}_{5}\left(m^{\alpha,\mu}\right)^{2}\biggr{)}\mathrm{d}t\mid X_{0}=x\Biggr{]}.\end{split}$ (6.6)
The explicit formula $v(x)=\Gamma_{2}x^{2}+\Gamma_{1}x+\Gamma_{0}$ can be derived as the solution to the Hamilton-Jacobi-Bellman equation where
$\Gamma_{2}$ $=\frac{-\beta+\sqrt{\beta^{2}+8\left(c_{1}+c_{3}+\tilde{c}_{1}\right)}}{4}$
$\Gamma_{1}$ $=-\frac{2\Gamma_{2}c_{3}c_{4}}{c_{1}\left(1-c_{2}\right)+\tilde{c}_{1}\left(1-\tilde{c}_{2}\right)^{2}+c_{3}+\tilde{c}_{5}}$
$\Gamma_{0}$ $=\frac{c_{1}c_{2}^{2}m^{2}+\left(\tilde{c}_{1}\tilde{c}_{2}^{2}+\tilde{c}_{5}\right)\left(m^{\alpha,\mu}\right)^{2}+\sigma^{2}\Gamma_{2}-\frac{1}{2}\Gamma_{1}^{2}+c_{3}c_{4}^{2}}{\beta}.$
@@ 6.2 Hyperparameters and Numerical Specifics
The results of the IH-MFCG-AC algorithm (Algorithm 2) are presented in figs. 17, 19 and 19. As expected, the learning of the global and local distributions reflects that of the optimal MFG distribution and the optimal MFC distribution, respectively. We observe that the global score is learned faster and with more accuracy than the local score, which is prone to outliers and instability. The optimal control is learned well within the support of the optimal distribution, but could possibly be expanded with a more advanced exploration strategy.
Figure 17: The histograms are the learned distributions generated using samples from the global score $\Sigma_{\varphi_{n}}$ (green) representing the global distribution $\hat{\mu}$ and the local score $\widetilde{\Sigma}_{\xi_{n}}$ (purple) representing the local score $\mu^{\alpha,\hat{\mu}}$ after $N=2\times 10^{6}$ iterations. The dashed line (blue) is the learned feedback control averaged over five runs with different initial samples. The benchmark solution to the MFCG is provided in orange. The $x$-axis shows the state variable $x$, the left $y$-axis refers to the value of the control $\alpha(x)$, and the right axis represents the probability density of $\mu(x)$.
Figure 18: The blue curve is a rolling average of the absolute error of the mean of samples produced from the global score function $\Sigma_{\varphi_{n}}$ (left)—denoted $\hat{m}_{t_{n}}$—and the local score function $\widetilde{\Sigma}_{\xi_{n}}$—denoted $m^{\alpha,\hat{\mu}}_{t_{n}}$—compared to the optimal mean $m$ from eq. 6.9. These values were averaged over five runs each with different random initial samples with the standard deviation given by the light blue shaded region. Large jumps are due to random outliers which result from the stochasticity of our algorithm.
Figure 19: The orange curve is the optimal value function for the MFCG problem. The blue dashed line is the learned value function given by the negative of critic $V_{\theta_{N}}$ averaged over five runs with different initial samples after $N=2\times 10^{6}$ iterations. Since the original optimization problem aims to minimize cost while our algorithm seeks to maximize reward, we take the negative of the critic function to make the problems equivalent. The light blue shaded region depicts one standard deviation from the learned value.